OpenStudy (anonymous):

You decide you will buy a stock only if it shows an overall increase over the next 30 days. The first 10 days it has an average daily increase of 0.30. The next 10 days it has an average daily decrease of 0.45. The last 10 days it has an average daily increase of 0.25. Will you buy the stock? Explain. 1 medal, one follow, and one testimonial PLEASE HELP

OpenStudy (anonymous):

@Christian_10_ do you think you could help me please?

OpenStudy (christian_10_):

Let's say the stock started out at price, X. So after 10days,the stock went up by(10days)times(30 cents a day)= 3.00. So after 10 days, the stock price is X+ 3. The next 10 days, the stock decreases by (10 days) times (45 cents a day) = -4.50.So the new price of the stock after 20days is ( X+3)− 4.50=X−1.50.So the price is below the starting price. Last 10 days the stock went up by(10days)times(25cents a day)= 2.5. So the final total is (X− 1.50)+2.5= X+$1 So the stock price went up one dollar over the whole month.

OpenStudy (anonymous):

@Christian_10_ thank you for your help. I understand. Your a life saver.! :-)

OpenStudy (christian_10_):

Your welcome! :) Thanks for the medal!